# Can Many-Shot In-Context Learning Help LLMs as Evaluators? A Preliminary Empirical Study

### v4

@@ Abstract
Utilizing Large Language Models (LLMs) as evaluators for evaluating the performance of LLMs has recently garnered attention. However, this kind of evaluation approach is affected by potential biases in LLMs, raising concerns about the accuracy and reliability of the evaluation results. To mitigate this issue, we propose and study two many-shot ICL prompts, which rely on two versions of many-shot ICL prompt templates for helping LLM evaluators to mitigate the potential biases in LLMs, Many-Shot with Reference (MSwR) and Many-Shot without Reference (MSoR). Concretely, the former utilizes in-context examples with model-generated rationales as guidance, and the latter without. Based on the designed prompts, we investigate the impact of scaling the number of in-context examples on the consistency and quality of the evaluation results. Experimental results show that advanced LLMs, such as GPT-4o, perform better in the many-shot regime than in the zero-shot regime. Furthermore, we reveal the symbol bias hidden in the selection bias of LLMs and propose a simple yet effective approach to mitigate the bias. Experimental results further verify the effectiveness of the symbol bias mitigation approach.
@@ A Preliminary Empirical Study
Mingyang Song, Mao Zheng, Xuan Luo Tencent MLPD nickmysong@tencent.com
@@ 2.2 Recalling Potential Biases in LLMs
To select the tested LLM and reveal the selection bias, we conducted a simple symbol selection experiment (set temperature=0.7 for all tested LLMs), using the designed prompt in Table 2 to let LLMs choose a symbol. The results are shown in Figure 4, and from these results, it can be observed that different LLMs tend different symbols. Meanwhile, we also evaluate the performance of LLaMA3-70B on the GSM8K dataset using Prompt(A) in Table 1 as shown in Figure 3. From the results, we can observe that the evaluation results of GPT-4o is relatively reasonable. Moreover, previous studies have demonstrated that GPT4 series models have the highest consistency with human evaluation.
@@ 2.3 Two Prompt Templates
In this section, we introduce two templates for LLM as an evaluator,many-shot with reference and many-shot without reference.
@@ 2.3.1 Many-Shot with Reference
Usually, when asking LLMs to select a score from 1-10 for evaluating answers to questions, providing reference answers may improve the evaluation quality of LLMs Wang et al. (2023b); Zheng et al. (2023). In the zero-shot ICL regime, no in-context examples are provided for the GPT-4o evaluator, which selects scores only depending on itself. In the few-shot ICL regime, a few in-context examples are provided for the GPT-4o evaluator. In the many-shot ICL regime, many in-context examples are provided for the GPT-4o evaluator. The differences among the three regimes above are as follows: (1) zero-shot, which lacks any reference information, causes LLM to score entirely based on its preferences. (2) few-shot, due to providing only a small number of examples, may lead to evaluation results that lack diversity. (3) many-shot, by offering a larger number of examples, can ensure diversity and quality in evaluations. Therefore, we present the prompt template MSwR, which uses model-generated rationales as the in-context examples. Specifically, MSoR-4 indicates the prompt template of MSoR using 4 shots with reference as a demonstration, as shown in Table 3.
@@ 2.3.2 Many-Shot without Reference
Previous studies Agarwal et al. (2024) find that in-context examples may limit the problem-solving approach of LLMs, so we propose a prompt template that does not need a reference for each in-context example. The designed prompt contains three parts: (1) a preamble, such as, “You will be provided questions similar to the ones below:”; (2) a list of unsolved inputs or problems.; (3) a few-shot prompt with outputs for the desired output format. Specifically, MSwR-(K=4)-4 indicates that using 4 shots without reference and 4 shots with reference, as shown in Table 4.
@@
Figure 5: Consistency between two versions of evaluation results. Concretely, the bar corresponding to "0" on the x-axis represents the number of samples with consistent and inconsistent ratings in comparing evaluation results obtained twice using GPT-4o as the evaluator in the zero-shot regime. In addition, the zero-shot generated rationales are used for MSwR and appended to the prompt for MSoR. The bar corresponding to "$2^{n}$" on the x-axis represents the consistency of using the GPT-4o as an evaluator in MSwR.
@@ 3.2 Consistency Evaluation
Figure 7: Results of random selection.
@@ 3.4 Revealing Symbol Bias
Inspired by our experiments, we swap the answers corresponding to symbols A and B, as shown in Figure 6. From Figure 8 Compare(A, B) and Compare(B, A), it can be seen that the results are different from the above experiments. Actually, the results of Compare(A, B) and Compare(A, B) should be similar, and the results of Compare(B, A) and Compare(B, A) should be similar. This phenomenon shows that symbol bias does exist when adopting GPT-4o as an evaluator.
@@ 6 Limitations
Considering the trade-off between costs and benefits, we do not verify too many in-context examples in the experiments, such as thousands of examples. Combining Figures 5 and Figure 9, it is not difficult to see that when the number of in-context examples increases to 256 and 512, although the consistency no longer improves, the evaluation quality has increased significantly. In addition, we consider that using GPT-4o as an evaluator in the many-shot regime is another evolution of the weak-to-strong strategy Burns et al. (2023), which uses many zero-shot evaluation results to generate a better one. As the long-context capabilities of LLMs improve, adding more in-context examples may reveal more valuable phenomena for studying LLMs as evaluators in the future.

### v5

@@ Abstract
Utilizing Large Language Models (LLMs) as evaluators to assess the performance of LLMs has garnered attention. However, this kind of evaluation approach is affected by potential biases within LLMs, raising concerns about the accuracy and reliability of the evaluation results of LLMs. To address this problem, we propose and study two many-shot In-Context Learning (ICL) prompt templates to help LLM evaluators mitigate potential biases: Many-Shot with Reference (MSwR) and Many-Shot without Reference (MSoR). Specifically, the former utilizes in-context examples with model-generated evaluation rationales as references, while the latter does not include these references. Using these prompt designs, we investigate the impact of increasing the number of in-context examples on the consistency and quality of the evaluation results. Experimental results show that advanced LLMs, such as GPT-4o, perform better in the many-shot regime than in the zero-shot and few-shot regimes. Furthermore, when using GPT-4o as an evaluator in the many-shot regime, adopting MSwR as the prompt template performs better than MSoR.
@@ A Preliminary Empirical Study
Mingyang Song, Mao Zheng, Xuan Luo, Yue Pan Tencent Hunyuan nickmysong@tencent.com
@@ 2.2 Recalling Potential Biases in LLMs
To select the tested LLM and reveal the selection bias, we conduct a simple symbol selection experiment (set temperature=0.7 for all tested LLMs), using the designed prompt in Table 2 to let LLMs randomly choose a symbol. The test results are shown in Figure 4, and from these results, it can be observed that different LLMs tend different symbols. Meanwhile, we also evaluate the performance of LLaMA3-70B on the GSM8K dataset using Prompt(A) in Table 1 as shown in Figure 3. From the results in Figure 4 and Figure 3, we observe that the results of GPT-4o are more reasonable than the others. Moreover, previous studies have demonstrated that GPT4 series models have the highest consistency with human evaluation.
@@ 2.3 Two Prompt Templates
In this section, we introduce two prompt templates for employing LLM as an evaluator: Many-Shot with Reference (MSwR) and Many-Shot without Reference (MSoR).
@@ 2.3.1 Many-Shot with Reference
Usually, when asking LLMs to select a score from 1-10 for evaluating answers to questions, providing reference answers may improve the evaluation quality of LLMs Wang et al. (2023b); Zheng et al. (2023). In the zero-shot ICL regime, no in-context examples are provided for the GPT-4o evaluator, which selects scores only depending on itself. In the few-shot ICL regime, a few in-context examples are provided for the GPT-4o evaluator. In the many-shot ICL regime, many in-context examples are provided for the GPT-4o evaluator. The differences among the three regimes above are as follows: (1) zero-shot, which lacks any reference information, causes LLM to score entirely based on its preferences. (2) few-shot, due to providing only a small number of examples, may lead to evaluation results that lack diversity. (3) many-shot, by offering a larger number of examples, can ensure diversity and quality in evaluations. Therefore, we present the prompt template MSwR, which uses model-generated rationales as the in-context examples. Specifically, MSwR-4 indicates the prompt template of MSwR using 4 shots with reference as a demonstration, as shown in Table 3.
@@ 2.3.2 Many-Shot without Reference
Previous studies Agarwal et al. (2024) find that in-context examples may limit the problem-solving approach of LLMs, so we propose a prompt template that does not need a reference for each in-context example. The designed prompt contains three parts: (1) a preamble, such as, “You will be provided questions similar to the ones below:”; (2) a list of unsolved inputs or problems.; (3) a few-shot prompt with outputs for the desired output format. Specifically, MSoR-(K=4)-4 indicates that using 4 shots without reference and 4 shots with reference, as shown in Table 4.
@@
Figure 5: Consistency between two versions of evaluation results. Concretely, the bar corresponding to "0" on the x-axis represents the number of samples with consistent and inconsistent ratings in comparing evaluation results obtained twice using GPT-4o as the evaluator in the zero-shot regime. In addition, the zero-shot generated rationales are used for MSwR and MSoR. The bar corresponding to "$2^{n}$" on the x-axis represents the consistency of using the GPT-4o as an evaluator in MSwR.
@@ 3.1 Experimental Settings
Figure 7: Results of randomly selecting different symbols.
3.2 Consistency Evaluation
@@ 3.4 Revealing Symbol Bias
Inspired by our experiments, we swap the answers corresponding to symbols A and B, as shown in Figure 6. From Figure 8 Compare(A, B) and Compare(B, A), it can be seen that the evaluation results are different from the experiments. Actually, the results of Compare(A, B) and Compare(A, B) should be similar, and the results of Compare(B, A) and Compare(B, A) should be similar. This phenomenon shows that symbol bias does exist when adopting GPT-4o as an evaluator.
@@ 6 Limitations
Considering the trade-off between costs and benefits, we do not verify too many in-context examples in the experiments, such as thousands of examples. Combining Figures 5 and Figure 9, it is not difficult to see that when the number of in-context examples increases to 256 and 512, although the consistency no longer improves, the evaluation quality has increased significantly. In addition, we consider that using GPT-4o as an evaluator in the many-shot regime is another evolution of the weak-to-strong strategy Burns et al. (2023), which uses many zero-shot evaluation results to generate a better one. As the long-context capabilities of LLMs improve, adding more in-context examples may reveal more valuable phenomena for studying LLMs as evaluators in the future. There are still many unexplored features of many-shot in-context learning. One possibility is that it can enable LLMs to simulate data constructed from shots, thereby increasing the diversity of synthetic data. Another possibility is to design benchmarks based on many-shot in-context examples and to assess the long-context LLMs based on the capabilities of understanding and imitating many shots.
Acknowledgments
We thank the three anonymous reviewers for carefully reading our paper and their insightful comments and suggestions.